Find the radius and centre of the circle:
(a) x^2 + y^2 - 2x + 4y + 1 = 0
1
Expert's answer
2021-01-26T01:52:01-0500

Given Equation :

"x^2+y^2-2x+4y+1=0"

which is the form of general equation of circle :

"x^2+y^2+2gx+2fy+c=0"

with center at "(-g,-f)" and radius being "\ r=\sqrt{g^2+f^2-c}"

In given equation,

"g=\frac{-2}{2}=-1\implies -g=1"

And "f=\frac{4}{2}=2\implies -f=-2"

And "c=1"

So, center of circle is "(-g,-f)\equiv(1,-2)"

And radius of circle will be "\ r=\sqrt{g^2+f^2-c}=\sqrt{(-1)^2+2^2-1}=\sqrt{4}=2" units.